OpenStudy (anonymous):

find a power model , by solving for y lny=3.2lnx+2.2 I'm lost with what's being asked, and the mathematical method to solve this

OpenStudy (apples):

Not really sure what a `power model' is, but a quick google leads me to believe it is a number of the form a * x^n. \[\log{y} = 3.2 \log{x} + 2.2\]\[e^{\log{y}} = e^{3.2 \log{x} + 2.2}\]\[y = (e^{\log{x}})^3.2 \cdot e^{2.2}\]\[y = e^{2.2} \cdot x^3.2\]

OpenStudy (anonymous):

ok, ya this seems pretty reasonable, referring back to previous problems i've completed. Thank you

OpenStudy (apples):

No problem :)